OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 25. What is the length of the altitude?

OpenStudy (anonymous):

Ahh got it!

OpenStudy (anonymous):

BD/AD = AB/AC CD/AD = AC/AB (inverse of above) so BD/AD = AD/CD

OpenStudy (anonymous):

BD*CD=AD^2 Now can you do it?

OpenStudy (anonymous):

@Sunshine447

OpenStudy (callisto):

@apple_pi I would do it in this way  a^2+b^2 = (16+25)^2 a^2 = c^2 + 16^2 -(1) b^2 = c^2 + 25^2 -(2) (1) + (2) a^2 + b^2 = c^2 + 16^2 + c^2 + 25^2 (16+25)^2 = 2c^2 + 16^2 + 25^2